# Dexterous Manipulation through Imitation Learning: A Survey

### v4

@@ V-C4 Dexterous and Bimanual Manipulation Datasets
Focusing on dexterous manipulation and hand–object interactions, ARCTIC [319] offers 2.1 million videos with precise 3D hand–object meshes and dynamic contact data, enabling the study of bimanual manipulation of articulated objects. DexGraspNet [320] provides 1.32 million grasps for 5,355 objects using ShadowHand, with each grasp physically validated in simulation to ensure stability. OAKINK2 [321] extends the scope to real-world bimanual object manipulation, comprising 627 sequences and 4.01 million frames from multi-view captures with detailed pose annotations for human bodies, hands, and objects. Tab. VIII presents a concise summary of the representative benchmarks and datasets discussed above.
From the perspective of dataset usability, several interdependent factors critically shape the practical value of an imitation learning dataset (as shown in Table IX). Foremost among these is sensor modality richness: datasets offering more comprehensive multi-modal signals (e.g., depth image, tactile feedback, proprioceptive states, force–torque readings, audio), generally deliver superior informational content and flexibility compared to RGB-only collections. Such multi-modal data facilitates more accurate perception, reasoning over physical contacts, and cross-modal policy learning. Annotation quality is another decisive factor, precise labels for object poses, manipulator or end effector positions, and task boundaries directly constrain the attainable performance of learned models. Task and scene diversity further influences generalization: datasets encompassing a wide spectrum of object categories, environmental configurations, and illumination conditions are better poised to support robust and transferable policies than those restricted to narrow, repetitive settings. For simulated or augmented datasets, physical realism is especially critical, as greater realism narrows the sim-to-real gap and boosts the likelihood that policies trained virtually will operate successfully on physical robots.
@@ VI Challenges and Future Directions in Imitation Learning-Based Dexterous Manipulation
Imitation learning-based dexterous manipulation poses unique challenges due to the inherent complexities of both IL and dexterous control. Despite significant advancements over the past decade, several challenges hinder its human-level dexterity and real-world applicability. Figure 6 summarizes the challenges based on research impact and technical difficulty, highlighting immediate priorities that are further discussed in the remainder of this section.
@@ VI-C Generalization to Novel Setups
Generalizing imitation learning-based dexterous manipulation policies is challenging due to task and environment variability, adaptive learning limitations, and cross-embodiment adaptability:
@@ VI-D1 Domain Randomization
One widely adopted strategy for bridging the sim-to-real gap is domain randomization. Instead of replicating real-world conditions exactly, domain randomization introduces variability in simulation parameters. This exposure helps policies generalize to unseen real-world scenarios [276]. However, domain randomization requires manual tuning of parameter distributions, which demands expert knowledge to balance realism and variability. It also struggles to address unmodeled dynamics, such as non-linear material deformations or high-frequency contact interactions, which are crucial for dexterous manipulation. Therefore, future research could focus on combining domain randomization with complementary sim-to-real techniques, such as real-world fine-tuning or adversarial domain adaptation, to achieve more robust and reliable transfer.
@@ VI-D3 Adversarial Domain Adaptation
Adversarial techniques have also emerged as a powerful way to minimize sim-to-real discrepancies. This approach has been applied to tasks where direct collection of real-world demonstrations is expensive [327]. In robotic manipulation, adversarial adaptation can be used in combination with IL by first imitating expert demonstrations in simulation and then fine-tuning the learned policy to fool a domain discriminator, ensuring smoother deployment in real-world settings. Future research could explore integrating adversarial adaptation with generative models or self-supervised representation learning to improve robustness and scalability, particularly for high-dimensional sensory inputs such as vision and tactile feedback.
@@ VI-D4 Hybrid Training and Online Adaptation
Beyond domain adaptation, hybrid training paradigms are gaining significant traction. In these approaches, policies are first pre-trained in simulation for rapid skill acquisition and subsequently fine-tuned on real-world data using IL, RL, or offline corrections [328]. This strategy combines the efficiency and scalability of simulation with the precision of real-world adaptation, thereby reducing the need for extensive real-world data collection. Additionally, self-supervised real-to-sim refinement, where real-world rollouts are used to iteratively adjust simulation parameters, can improve the fidelity of simulated environments and facilitate bi-directional transfer.
@@ VI-F Safety, Robustness, and Social Compliance
Ensuring safety, robustness, and social compliance is crucial for real-world dexterous robotics, requiring risk prevention, adaptive error recovery, and human-aware behavior for seamless integration.
@@
Real-world dexterous manipulation demands reliable error detection, recovery, and adaptability in dynamic, unstructured environments. Detecting failures is challenging due to sensor noise, occlusions, and unpredictable interactions, while recovery strategies like re-grasping or trajectory replanning must be executed in real time to maintain stability and task continuity. Future research should address two key aspects. First, large-scale failure datasets and standardized benchmarks are essential for improving data-driven recovery policies. Establishing comprehensive datasets and evaluation protocols for failure detection, uncertainty estimation, and recovery effectiveness would provide a foundation for training and benchmarking robust policies. Second, self-supervised learning for multi-modal anomaly detection could enable robots to autonomously refine their error detection capabilities. By leveraging visual, tactile, and proprioceptive feedback, robots could learn to recognize and anticipate failures in real time, improving adaptability and robustness in dynamic environments.
@@
Safety is equally critical, particularly in real-world deployments where unpredictable interactions and dynamic conditions pose significant risks. In dexterous manipulation, safety considerations involve collision avoidance, force regulation, and compliance control, particularly when interacting with fragile objects or operating near humans. However, achieving these safety measures requires handling varying contact conditions, but sensor noise, occlusions, and data processing delays can reduce reliability. Additionally, while compliant actuators and soft robotic designs help mitigate impact forces, integrating these hardware safety mechanisms involves trade-offs between control precision, responsiveness, and durability.
@@
To address safety constraints in IL for real-world dexterous manipulation, recent studies have investigated constrained policy optimization [329] and Lyapunov-based safe learning [330]. Constrained policy optimization integrates hard limits on torque, contact force, or joint velocity directly into the training objective, often via Lagrangian relaxation or primal-dual methods. For instance, Constrained Trust Region Policy Optimization (C-TRPO) [331] augments the trust-region update rule by incorporating actuator and contact-force limits, ensuring that each policy update remains close to expert demonstrations while strictly respecting physical safety bounds. In parallel, Lyapunov-based approaches synthesize control policies that provably satisfy safety specifications, constructing Lyapunov functions as certificates of stability. In dexterous manipulation, Neural Lyapunov Control (NLC) [332] has been adapted to learn neural network policies that maintain force closure and avoid excessive contact forces. Recent work combines these paradigms. Xiao et al. [333] propose SafeDiffuser, which augments diffusion-policy-based planning with control barrier functions to enforce real-time safety constraints. By conditioning the diffusion model on expert demonstrations and integrating CBF-based corrections at every denoising step, the method achieves robust generalization to unseen scenarios while guaranteeing safety-critical bounds.
@@
Beyond technical safety, social compliance is essential for real-world deployment yet remains underexplored, particularly in HRI settings. In the context of dexterous manipulation, prior research has investigated various approaches, such as RL-based methods (e.g., preference-based RL[334] and inverse RL[335]), MPC frameworks (e.g., socially aware Model Predictive Path Integral control[336]), and game-theoretic models (e.g., Nash equilibrium formulations[337]). However, achieving social compliance remains challenging due to the inherent ambiguity of human comfort, spatial preferences, and social norms, which are often context-dependent and difficult to model explicitly. Additionally, soft social constraints, such as maintaining legibility of motion, avoiding intrusive actions, or signaling intent, are subjective and require learning from diverse human feedback, which is costly and time-consuming. Another challenge lies in the lack of standardized evaluation metrics and benchmarking protocols to assess social compliance. Existing metrics, such as trajectory smoothness, proximity compliance, or motion legibility, are often evaluated qualitatively or through small-scale user studies without a consistent framework. Moreover, there is a lack of publicly available datasets and simulation environments specifically designed for evaluating social compliance in dexterous manipulation, unlike traditional benchmarks for grasping or motion planning (e.g., RLBench [338]or ManiSkill3 [339]). Without well-defined metrics, datasets, and test scenarios, progress in this area remains fragmented and difficult to measure systematically.
@@ VI-G High-Precision and Micro-Scale Manipulation
While the above discussion focuses on macro-scale object manipulation (e.g., bottles, tools), scaling down to the millimeter or micrometer level (e.g., electronic components, micro-machines) introduces fundamentally different challenges. At these scales, perception limitations, such as low sensor resolution, visual noise, and occlusion, become critical. Actuation requirements also rise sharply, demanding ultra-precise force and motion control within micrometer tolerances to prevent damage to delicate objects. Conventional actuators frequently suffer from backlash, hysteresis, and limited resolution, while achieving high control precision becomes a bottleneck. These challenges have driven the development of advanced control methods [340, 341, 342], including force-feedback, adaptive impedance, and hybrid position-force strategies designed specifically for micro-scale tasks.
Moreover, data collection at micro-scales remains a significant challenge. Acquiring large, diverse, and high-quality demonstrations is both costly and time-intensive, given the specialized equipment and delicate experimental setups required. Synthetic data generation through high-fidelity physics-based simulation is a promising direction, but it requires accurate modeling of micro-scale physical effects, such as adhesion forces, surface roughness, or fluid dynamics, which are often negligible in macro-scale tasks. From an algorithmic perspective, recent research [74] underscores the potential of hierarchical IL frameworks in addressing dexterous manipulation in micro-scale tasks. These techniques, together with advances in small-scale robots[343], present a promising pathway for enabling autonomous micro-manipulation tasks, such as precision assembly and surgical procedures.
@@ VII Conclusion
The future of dexterous manipulation holds great potential, with applications ranging from industrial automation to healthcare and service robotics. By continuing to push the boundaries of IL and robotic manipulation, researchers and practitioners can pave the way for more capable, adaptable, and intelligent robotic systems. These advancements will not only improve the efficiency and safety of robotic tasks but also open up new possibilities for human-robot collaboration and interaction.
@@ Acknowledgment
We would like to thank Huapeng Li, Qianyi Wang, Shouzheng Wang, and Mingwu Liu for their contributions to the completion of this survey.

### v5

@@ V-C4 Dexterous and Bimanual Manipulation Datasets
Focusing on dexterous manipulation and hand–object interactions, ARCTIC [fan2023arctic] offered 2.1 million videos with precise 3D hand–object meshes and dynamic contact data, enabling the study of bimanual manipulation of articulated objects. DexGraspNet [wang2023dexgraspnet] provided 1.32 million grasps for 5,355 objects using ShadowHand, with each grasp physically validated in simulation to ensure stability. OAKINK2 [2403.19417] extended the scope to real-world bimanual object manipulation, comprising 627 sequences and 4.01 million frames from multi-view captures with detailed pose annotations for human bodies, hands, and objects. Table VIII presents a concise summary of the representative benchmarks and datasets discussed above.
From the perspective of dataset usability, several interdependent factors critically shape the practical value of an imitation learning dataset (as shown in Table IX). Foremost among these is sensor modality richness: datasets offering more comprehensive multi-modal signals (e.g., depth image, tactile feedback, proprioceptive states, force–torque readings, audio) generally deliver superior informational content and flexibility compared to RGB-only collections. Such multi-modal data facilitates more accurate perception, reasoning over physical contacts, and cross-modal policy learning. Annotation quality is another decisive factor; precise labels for object poses, manipulator or end effector positions, and task boundaries directly constrain the attainable performance of learned models. Task and scene diversity further influences generalization: datasets encompassing a wide spectrum of object categories, environmental configurations, and illumination conditions are better poised to support robust and transferable policies than those restricted to narrow, repetitive settings. For simulated or augmented datasets, physical realism is especially critical, as greater realism narrows the sim-to-real gap and boosts the likelihood that policies trained virtually will operate successfully on physical robots.
@@ VI Challenges and Future Directions in Imitation Learning-Based Dexterous Manipulation
IL-based dexterous manipulation poses unique challenges due to the inherent complexities of both IL and dexterous control. Despite significant advancements over the past decade, several challenges hinder its human-level dexterity and real-world applicability. Fig. 6 summarizes the challenges based on research impact and technical difficulty, highlighting immediate priorities further discussed in this section.
@@ VI-C Generalization to Novel Setups
Generalizing IL-based dexterous manipulation policies is challenging due to task and environment variability, adaptive learning limitations, and cross-embodiment adaptability:
@@ VI-D1 Domain Randomization
One widely adopted strategy for bridging the sim-to-real gap is domain randomization. Instead of replicating real-world conditions exactly, domain randomization introduces variability in simulation parameters. This exposure helps policies generalize to unseen real-world scenarios [tobin2017domain]. However, domain randomization requires manual tuning of parameter distributions, which demands expert knowledge to balance realism and variability. It also struggles to address unmodeled dynamics, such as non-linear material deformations or high-frequency contact interactions, which are crucial for dexterous manipulation. Therefore, future research could focus on combining domain randomization with complementary sim-to-real techniques, such as real-world fine-tuning or adversarial domain adaptation, to achieve more robust and reliable transfer.
@@ VI-D3 Adversarial Domain Adaptation
Adversarial techniques have also emerged as a powerful way to minimize sim-to-real discrepancies. This approach has been applied to tasks where direct collection of real-world demonstrations is expensive [tzeng2017adversarial]. In robotic manipulation, adversarial adaptation can be used in combination with IL by first imitating expert demonstrations in simulation and then fine-tuning the learned policy to fool a domain discriminator, ensuring smoother deployment in real-world settings. Future research could explore integrating adversarial adaptation with generative models or self-supervised representation learning to improve robustness and scalability, particularly for high-dimensional sensory inputs such as vision and tactile feedback.
@@ VI-D4 Hybrid Training and Online Adaptation
Beyond domain adaptation, hybrid training paradigms are gaining significant traction. In these approaches, policies are first pre-trained in simulation for rapid skill acquisition and subsequently fine-tuned on real-world data using IL, RL, or offline corrections [rajeswaran2018learning]. This strategy combines the efficiency and scalability of simulation with the precision of real-world adaptation, thereby reducing the need for extensive real-world data collection. Additionally, self-supervised real-to-sim refinement, where real-world rollouts are used to iteratively adjust simulation parameters, can improve the fidelity of simulated environments and facilitate bi-directional transfer.
@@ VI-F Safety, Robustness, and Social Compliance
Ensuring safety, robustness, and social compliance is crucial for real-world dexterous robotics, requiring risk prevention, adaptive error recovery, and human-aware behavior for seamless integration. Real-world dexterous manipulation demands reliable error detection, recovery, and adaptability in dynamic environments. Detecting failures is challenging due to sensor noise, occlusions, and unpredictable interactions, while recovery strategies like re-grasping or trajectory replanning must be executed in real time to maintain stability and task continuity. Future research should address two key aspects. First, large-scale failure datasets and standardized benchmarks are essential for improving data-driven recovery policies. Establishing comprehensive datasets and evaluation protocols for failure detection, uncertainty estimation, and recovery effectiveness would provide a foundation for training and benchmarking robust policies. Second, self-supervised learning for multi-modal anomaly detection could enable robots to autonomously refine their error detection capabilities. By leveraging visual, tactile, and proprioceptive feedback, robots could learn to recognize and anticipate failures in real time, improving adaptability and robustness in dynamic environments.
@@
Safety is equally critical, particularly in real-world deployments where unpredictable interactions and dynamic conditions pose significant risks. In dexterous manipulation, safety considerations involve collision avoidance, force regulation, and compliance control, particularly when interacting with fragile objects or operating near humans. However, achieving these safety measures requires handling varying contact conditions, but sensor noise, occlusions, and data processing delays can reduce reliability. Additionally, while compliant actuators and soft robotic designs help mitigate impact forces, integrating these hardware safety mechanisms involves trade-offs between control precision, responsiveness, and durability. Different paradigms also entail distinct safety concerns. End-to-end learning directly maps perception to control, but its black-box nature complicates safety verification and error tracing. Motion module learning decomposes the pipeline into structured stages, improving generalization and safety at the cost of intermediate complexity.
@@
To address safety constraints in IL for real-world dexterous manipulation, recent studies have investigated constrained policy optimization [achiam2017constrained] and Lyapunov-based safe learning [chow2018lyapunov]. Constrained policy optimization integrates hard limits on torque, contact force, or joint velocity directly into the training objective, often via Lagrangian relaxation or primal-dual methods. For instance, Constrained Trust Region Policy Optimization (C-TRPO) [milosevic2024embedding] augmented the trust-region update rule by incorporating actuator and contact-force limits, ensuring that each policy update remains close to expert demonstrations while strictly respecting physical safety bounds. In parallel, Lyapunov-based approaches synthesized control policies that provably satisfy safety specifications, constructing Lyapunov functions as certificates of stability [khansari2011learning, xu2020learning, xu2021learning]. In dexterous manipulation, Neural Lyapunov Control (NLC) [chang2019neural] has been adapted to learn neural network policies that maintain force closure and avoid excessive contact forces. Recent work combined these paradigms. SafeDiffuser [xiao2023safediffuser] augmented diffusion-policy-based planning with control barrier functions to enforce real-time safety constraints. By conditioning the diffusion model on expert demonstrations and integrating CBF-based corrections at every denoising step, the method achieves robust generalization to unseen scenarios while guaranteeing safety-critical bounds.
@@
Beyond technical safety, social compliance is essential for real-world deployment yet remains underexplored, particularly in HRI settings. In the context of dexterous manipulation, prior research has investigated various approaches, such as RL-based methods (e.g. preference-based RL [an2023direct] and inverse RL [das2021model]), MPC frameworks (e.g. socially aware Model Predictive Path Integral control [trevisan2024biased]), and game-theoretic models (e.g. Nash equilibrium formulations [bansal2020bayesian]). However, achieving social compliance remains challenging due to the inherent ambiguity of human comfort, spatial preferences, and social norms, which are often context-dependent and difficult to model explicitly. Additionally, soft social constraints, such as maintaining legibility of motion, avoiding intrusive actions, or signaling intent, are subjective and require learning from diverse human feedback, which is costly and time-consuming. Another challenge lies in the lack of standardized evaluation metrics and benchmarking protocols to assess social compliance. Existing metrics, such as trajectory smoothness, proximity compliance, or motion legibility, are often evaluated qualitatively or through small-scale user studies without a consistent framework. Moreover, there is a lack of publicly available datasets and simulation environments specifically designed for evaluating social compliance in dexterous manipulation, unlike traditional benchmarks for grasping or motion planning (e.g., RLBench [james2019rlbench] or ManiSkill3 [taomaniskill3]). Without well-defined metrics, datasets, and test scenarios, progress in this area remains fragmented and difficult to measure systematically.
@@ VI-G High-Precision and Micro-Scale Manipulation
While the above discussion focuses on macro-scale object manipulation (e.g., bottles, tools), scaling down to the millimeter or micrometer level (e.g., electronic components, micro-machines) introduces fundamentally different challenges. At these scales, perception limitations, such as low sensor resolution, visual noise, and occlusion, become critical. Actuation requirements also rise sharply, demanding ultra-precise force and motion control within micrometer tolerances to prevent damage to delicate objects. Conventional actuators frequently suffer from backlash, hysteresis, and limited resolution, while achieving high control precision becomes a bottleneck. These challenges have driven the development of advanced control methods [zhong2023spatial, xin2025dynamic, wang2024data], including force-feedback, adaptive impedance, and hybrid position-force strategies designed specifically for micro-scale tasks.
Moreover, data collection at micro-scales remains a significant challenge. Acquiring large, diverse, and high-quality demonstrations is both costly and time-intensive, given the specialized equipment and delicate experimental setups required. Synthetic data generation through high-fidelity physics-based simulation is a promising direction, but it requires accurate modeling of micro-scale physical effects, such as adhesion forces, surface roughness, or fluid dynamics, which are often negligible in macro-scale tasks. From an algorithmic perspective, recent research [kim2025srt] underscored the potential of hierarchical IL frameworks in addressing dexterous manipulation in micro-scale tasks. These techniques, together with advances in small-scale robots[dong2022untethered], presented a promising pathway for enabling autonomous micro-manipulation tasks, such as precision assembly and surgical procedures.
VI-H Roadmap for IL-based Dexterous Manipulation
To summarize, we propose a potential roadmap to address these issues, distinguishing between low-hanging fruits and grand, long-term challenges.
VI-H1 Low-Hanging Fruits
Data collection and benchmarking are high-impact areas that require relatively low research difficulty. Efforts focused on synthetic data generation, domain randomization, and crowdsourced teleoperation can quickly scale up the diversity of training datasets, which is essential for improving the generalization capabilities of models. Furthermore, establishing standardized benchmarking frameworks and evaluation protocols will streamline the process of comparing performance across different studies. These actions will enable faster progress in dexterous manipulation and contribute to broader adoption of the proposed techniques.
VI-H2 Grand, Long-Term Challenges
Sim-to-real transfer and real-time control are high-impact challenges that involve significant research difficulty. Bridging the sim-to-real gap will require breakthroughs in domain adaptation, hybrid training paradigms, and better computational resources. These advances are crucial for ensuring the successful deployment of models trained in simulation into real-world environments. Real-time control, particularly for high-dimensional tasks like dexterous manipulation, presents a need for more efficient algorithms, better control strategies, and seamless hardware integration. Solving these challenges will be essential for creating practical, high-performance systems capable of operating in dynamic, unstructured environments.
VI-H3 Safety and High-Precision Manipulation
While safety and high-precision manipulation are specialized concerns, they remain essential for real-world applications. These areas are high-difficulty but relatively lower in impact compared to sim-to-real transfer and real-time control. However, safety remains critical for human-robot interaction, especially in unstructured environments, and high-precision control is crucial for tasks like micro-manipulation. Future research should continue to address these challenges for real-world deployment.
@@ VII Conclusion
The future of dexterous manipulation holds great potential, with applications ranging from industrial automation to healthcare and service robotics. By continuing to push the boundaries of IL and robotic manipulation, researchers and practitioners can pave the way for more capable, adaptable, and intelligent robots. These advances will not only improve the efficiency and safety of robotic tasks but also open up new possibilities for human-robot collaboration and interaction.
@@ Acknowledgment
We would like to thank Huapeng Li, Qianyi Wang, Mingwu Liu, Shouzheng Wang and Zijing Yang for their contributions to the completion of this survey.
[Uncaptioned image] Shan An (Senior Member, IEEE) received the B.E. degree from Tianjin University in 2007, the M.E. degree from Shandong University in 2010, and the Ph.D. degree in computer science and engineering from Beihang University in 2022. He is currently an Associate Professor with the School of Electrical and Information Engineering, Tianjin University. His research interests include dexterous manipulation, extended reality for robotics, and robotic vision. Prior to his academic career, he accumulated 14 years of extensive industrial experience at the China Academy of Space Technology, Alibaba, and JD Group. He was elected Fellow of the Institution of Engineering and Technology (IET) in 2023 and Fellow of the British Computer Society (BCS) in 2024.
[Uncaptioned image] Ziyu Meng received the B.S. degree from Shandong University, Jinan, China, in 2022. He is currently pursuing the Ph.D. degree in pattern recognition and intelligent systems at Shandong University, Jinan, China. His current research interests include whole-body control for humanoid robots.
[Uncaptioned image] Chao Tang is currently a postdoctoral researcher at the Division of Robotics, Perception and Learning, KTH Royal Institute of Technology. He received his Ph.D. degree from Southern University of Science and Technology in 2025 and his M.S. degree from the Georgia Institute of Technology in 2020. His research interests include robotic manipulation and grasping, mobile manipulation, human–robot interaction, and semantic reasoning for robots.
[Uncaptioned image] Yuning Zhou received the B.E. degree in Mechanical Engineering from the University of Leeds and Southwest Jiaotong University (dual-degree program) in 2022, and the M.Sc. degree in Robotics, Systems and Control from ETH Zurich in 2025. His research interests include the design, simulation, and benchmarking of anthropomorphic robotic hands.
[Uncaptioned image] Tengyu Liu received the B.S. degree in computer science from the University of Illinois Urbana-Champaign (UIUC), Champaign, IL, USA, and the M.S. and Ph.D. degrees in computer science from the University of California, Los Angeles (UCLA), Los Angeles, CA, USA, in 2021. He is currently a Senior Research Scientist at Beijing Institute of General Artificial Intelligence (BIGAI), Beijing, China. His research interests lie at the intersection of 3D computer vision, computer graphics, and robotics, specifically focusing on generalizable dexterous grasping, manipulation, and whole-body control of humanoid and quadruped robots.
[Uncaptioned image] Fangqiang Ding is a Postdoctoral Associate with the Department of Mechanical Engineering at the Massachusetts Institute of Technology (MIT), Cambridge, MA, USA, working with Dr. Hermano Igo Krebs in The 77 Lab. He was previously a Postdoctoral Fellow at Technion – Israel Institute of Technology, working with Dr. Or Litany. He received the Ph.D. degree in Robotics and Autonomous Systems from the School of Informatics, University of Edinburgh, U.K., and the B.E. degree from Tongji University, Shanghai, China. He was selected as an RSS Pioneer in 2025 for his work on robust spatial perception with 4D radar for mobile autonomy. My research centers on reliable and affordable Physical AI, enabling AI-integrated physical systems (e.g., autonomous vehicles, robots, and IoT) to operate responsibly around humans and deliver societal benefits at scale in the physical world.
[Uncaptioned image] Yao Mu is an Assistant Professor at the Institute of Artificial Intelligence, Shanghai Jiao Tong University. He received his Ph.D. from the University of Hong Kong in 2025 and his Master’s from Tsinghua University in 2021. His research in embodied intelligence, reinforcement learning, robot control, and autonomous driving has produced over 40 papers in top venues including RSS, NeurIPS, ICML, ICLR, and CVPR, accumulating more than 2500 citations. Dr. Yao’s work has earned notable recognition including the ECCV Embodied Intelligence Workshop Best Paper Award, IEEE ICCAS 2020 Best Student Paper Award, and IEEE IV2021 Best Student Paper Nomination. He is a recipient of the Hong Kong Ph.D. Fellowship, University of Hong Kong Presidential Scholarship, and the National Scholarship of China. His research advances AI systems capable of effectively interacting with the physical world.
[Uncaptioned image] Shufang Zhang is currently an associate professor in the School of Electrical and Information Engineering, Tianjin University, Tianjin, China. She received her M.S. and Ph.D. degrees from Tianjin University in 2004 and 2007, respectively. Her research interests include Robot and SLAM, Artificial Intelligence, and Embodied Intelligence.
[Uncaptioned image] Ran Song (Senior Member, IEEE) is a Professor with the School of Control Science and Engineering, Shandong University, China since 2020. Before his current post, he was a senior lecturer at the University of Brighton, UK. He received his Ph.D. degree in electronic engineering from the University of York, UK in 2009 and his first degree from Shandong University in 2005. He has published more than 100 papers in peer-reviewed international conference proceedings and journals. His research interests lie in 3D visual perception, 3D vision for robotics, and robot learning.
[Uncaptioned image] Wei Zhang (Senior Member, IEEE) received the Ph.D. degree in electronic engineering from the Chinese University of Hong Kong in 2010. He is currently a Professor with the School of Control Science and Engineering, Shandong University, Jinan, China. His research interests include computer vision and robotics. He has served as a program committee member and a reviewer for various international conferences and journals.
[Uncaptioned image] Zeng-Guang Hou (Fellow, IEEE) received the B.E. and M.E. degrees in electrical engineering from Yanshan University (formerly Northeast Heavy Machinery Institute), Qinhuangdao, China, in 1991 and 1993, respectively, and the Ph.D. degree in electrical engineering from the Beijing Institute of Technology, Beijing, China, in 1997. He is currently a Professor with the State Key Laboratory of Multimodal Artificial Intelligence Systems, Institute of Automation, Chinese Academy of Sciences. His research interests include neural networks, robotics, and intelligent systems. He is serving as a VP of the Asia Pacific Neural Network Society (APNNS) and Chinese Association of Automation (CAA). He is an associate editor of IEEE Transactions on Neural Networks and Learning Systems, IEEE Transactions on Cognitive and Developmental Systems, and Neural Networks, etc. He is on the Board of Governors of International Neural Network Society (INNS). He was the Chair of Neural Network Technical Committee (NNTC) of Computational Intelligence Society (CIS), IEEE. Dr. Hou was a recipient of Neural Networks Best Paper Award in 2022, IEEE Transactions on Neural Networks Outstanding Paper Award in 2013.
[Uncaptioned image] Hong Zhang (Life Fellow, IEEE) received his Ph.D. in Electrical Engineering from Purdue University in 1986. He was a Professor in the Department of Computing Science, University of Alberta, Canada, for over 30 years before he joined the Southern University of Science and Technology (SUSTech), China, in 2020, where he is currently a Chair Professor. Dr. Zhang served as the Editor-in-Chief of IROS Conference Paper Review Board (2020-2022) and as a member of the IEEE Robotics and Automation Society Administrative Committee (2023-25). He is a Life Fellow of IEEE and a Fellow of the Canadian Academy of Engineering. His research interests include robotics, computer vision, and image processing.